VIBRATIONAL SPECTROSCOPY OF SYMPATHETICALLY COOLED CaH^{+} MOLECULAR IONS

\underline{NCAMISO B. KHANYILE}, JAMES E. GOEDERS and KENNETH R. BROWN, Department of Chemistry, Georgia Institute of Technology, Atlanta, GA 30332.

The search for time variation in the fundamental constants of nature such as the fine structure constant(α) and the proton/electron mass ratio(μ), is an area of active research^{a}.Comparing the vibrational overtones of CaH^{+} with electronic transitions in atoms has been proposed as a means to detect possible time variation of μ ^{b}. Before these precision measurements can be realized, the survey spectroscopy needs to be performed. We describe our experiments using a Coulomb crystal of sympathetically cooled CaH^{+} and laser-cooled Ca^{+} ions to measure the vibrational overtones by resonance-enhanced multiphoton photo-dissociation (REMPD) in a linear Paul trap. The dissociation of CaH^{+} is detected by observing the change in the crystal composition by monitoring the Ca^{+} fluorescence. Future single ion experiments for the precision measurement are also discussed.

aJ. Uzan, Rev. Mod. Phys. 75, 403 (2003).

bM. Kajita and Y. Moriwaki, J. Phys. B: At. Mol. Opt. Phys. 42, 154022(2009).